### v1

# SCAN: Semantic Document Layout Analysis for Textual and Visual Retrieval-Augmented Generation

Abstract
With the increasing adoption of Large Language Models (LLMs) and Vision-Language Models (VLMs), rich document analysis technologies for applications like Retrieval-Augmented Generation (RAG) and visual RAG are gaining significant attention. Recent research indicates that using VLMs can achieve better RAG performance, but processing rich documents still remains a challenge since a single page contains large amounts of information. In this paper, we present SCAN (SemantiC Document Layout ANalysis), a novel approach enhancing both textual and visual Retrieval-Augmented Generation (RAG) systems working with visually rich documents. It is a VLM-friendly approach that identifies document components with appropriate semantic granularity, balancing context preservation with processing efficiency. SCAN uses a coarse-grained semantic approach that divides documents into coherent regions covering continuous components. We trained the SCAN model by fine-tuning object detection models with sophisticated annotation datasets. Our experimental results across English and Japanese datasets demonstrate that applying SCAN improves end-to-end textual RAG performance by up to 9.0% and visual RAG performance by up to 6.4%, outperforming conventional approaches and even commercial document processing solutions.
SCAN: Semantic Document Layout Analysis for Textual and Visual Retrieval-Augmented Generation
Yuyang Dongthanks: Equal contribution, Nobuhiro Ueda11footnotemark: 1, Krisztián Boros, Daiki Ito, Takuya Sera, Masafumi Oyamada NEC Corporation {dongyuyang, nobuhiro-ueda, krisztian-boros, ito-daiki, takuya-sera, oyamada}@nec.com
1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2021); Gao et al. (2024); Fan et al. (2024) technology enables Large Language Models (LLMs) to provide more accurate responses to user queries by retrieving and leveraging relevant knowledge and documents. These knowledge sources such as company financial reports, web pages, insurance manuals, and academic papers often contain complex charts, tables, diagrams, and other non-textual elements, collectively referred to as rich documents. Enabling RAG to accurately support these rich documents is an important research problem.
RAG frameworks are generally categorized into textual RAG and visual RAG. For textual RAG with rich documents, we need to convert documents (PDFs, images, etc.) into text, use a text retrieval, and then generate responses by LLMs. In contrast, visual RAG can directly process rich documents as images using multimodal embedding and retrieval, and then uses Vision Language Models (VLMs) to generate responses.
Figure 1: Conventional fine-grained layout analysis result (Left, DocLayout-YOLO) vs. Our coarse-grained semantic layout analysis result (Right, SCAN).
Figure 2: Overview of applying our SCAN model to current textual and visual RAG systems.
1.1 Challenges
The key part for textual RAG with rich documents is to convert documents into textual formats such as markdown. Recent research Zhang et al. (2025); Fu et al. (2024) indicates that using VLMs for document conversion yields better results than traditional OCR methods. However, it is still a challenge to have a VLM perform text conversion on rich document pages, since a whole page can contain a large amount of complicated content, which can lead to incomplete results and conversion errors.
On the other hand, visual RAG directly retrieves the images of rich document pages, but faces a similar challenge when using a VLM to process visual question answering (VQA) on those rich document pages. This is because user queries often target only specific portions of a document page, meaning that other parts of this page are noise, leading to low accuracy.
Therefore, the common challenge is to use a VLM to process (text conversion or VQA) an entire rich document page containing abundant information at once. One potential solution is to further divide a document page into small regions. Traditional document layout analysis technologies such as DocLayout-YOLO Zhao et al. (2024b) can achieve this objective, but they focus on fine-grained analysis, breaking down content into small components such as titles, paragraphs, tables, figures, and captions. This approach could lose important context when processing isolated components and potentially lead to reduced RAG accuracy. In our experiments, conventional layout analysis methods with VLM text conversion and VQA reduce accuracy by 5.8%–8.7% for textual RAG and 28.7%–40.5% for visual RAG, respectively.
1.2 Contributions
To address these challenges, we propose SCAN, a novel approach that performs VLM-friendly semantic document layout analysis with “coarse granularity.” Figure 1 compares the layout analysis results between SCAN and a conventional fine-grained method. SCAN can semantically divide regions into boxes that cover continuous components related to the same topic. For example, each of the semantic_box [3], [4], and [5] corresponds to independent topics of IT Services, Social Infrastructure, and Others.
To train a powerful SCAN model, we annotated more than 24k document pages with semantic layout labels. SCAN model is fine-tuned from pre-trained object detection models with this training data. We also designed a specific box selection strategy and post-processing techniques for RAG applications. Figure 2 gives an overview of applying our SCAN model into both textual and visual RAG pipelines. Concretely, each page of an input document is treated as an image and decomposed into semantic chunks by our SCAN model. For textual RAG, the images of semantic chunks are passed to a VLM that performs text conversion, and then input into existing textual RAG systems. For visual RAG, the resulting image chunks can be the retrieval targets that are directly input to existing visual RAG systems.
We test SCAN’s performance using three datasets featuring both English and Japanese documents. Although SCAN is trained on Japanese data, the experiments show that it can achieve good performance on English benchmarks. Experimental results show that in textual RAG, applying SCAN can improve end-to-end performance by 1.7%–9.0%, while in visual RAG, SCAN can enhance end-to-end performance by 3.5%–6.4%. Moreover, although our SCAN requires multiple processes on VLMs rather than a single page-level processing, the total computational time is roughly the same while the total token usage is reduced.
2 Related Work
2.1 Document Layout Analysis
CNN backbones such as DocLayout-YOLO (Zhao et al., 2024b) and transformer backbones such as DiT (Li et al., 2022), LayoutLMv3 (Huang et al., 2022), LayoutDETR (Yu et al., 2023), Beehive (Auer et al., 2024) have been proposed for high-performance document layout analysis which are trained with synthetic and human-annotated datasets Zhong et al. (2019); Li et al. (2020); Pfitzmann et al. (2022). End-to-end document conversion systems have been developed using the above models. Docling (Auer et al., 2024), Marker (Datalab, 2024), and MinerU (Wang et al., 2024) provide a comprehensive pipeline for document layout analysis and text conversions. There are also production systems such as Azure Document Intelligence microsoft (2025) and LlamaParse Premium LlamaParse.AI (2025).
2.2 VLMs for Document Conversion
Vision-Language Models (VLMs) have emerged as powerful tools for multimodal understanding tasks. Open models such as the Qwen-VL series Bai et al. (2025), or InternVL Chen et al. (2024b) have also demonstrated impressive capabilities in visual document understanding. There are also small-sized OCR-specialized VLMs such as GOT Wei et al. (2024), Nougat Blecher et al. (2023), DocVLM Nacson et al. (2024), and olmOCR Poznanski et al. (2025).
2.3 Textual and visual RAG
The rapid progress of LLMs has further strengthened RAG, enabling models to inject external knowledge into responses with higher precision and coverage Lewis et al. (2021); Gao et al. (2024); Fan et al. (2024). Typical RAG pipelines in previous works first converted document images or PDFs into plain text, and only then performed indexing and retrieval over the extracted passages Zhang et al. (2025). Recent results show that using VLMs for document text conversion yields better results than traditional OCR tools Zhang et al. (2025); Fu et al. (2024). On the other hand, with the increasing availability of multimodal embedding and VLMs, there is a growing interest in multimodal RAG systems that can directly index and retrieve images, and can leverage VLMs for answer generation Yu et al. (2025); Tanaka et al. (2025); Faysse et al. (2024).
3 Method
3.1 SCAN Model Training
Given a single page image from a rich document, the semantic document layout analysis problem is to identify and delineate different regions of the image using bounding boxes, each associated with a specific semantic type. This formulation follows classical layout analysis approaches and can be viewed as a specialized object detection task. Accordingly, we fine-tune pre-trained object detection models on a dataset with semantic layout labels, leveraging their robust feature extraction capabilities while adapting them to the requirements of document layout analysis.
3.1.1 Training Data
We define the semantic boxes are regions that contain content unified by a single topic or meaning. Unlike structural divisions (e.g., paragraphs or sections), semantic boxes are defined by their coherent meaning. In other words, all content within a box relates to the same subject and can be understood independently, without needing context from surrounding boxes. A semantic box might contain various structural elements, multiple paragraphs, sections, tables, or charts, as long as they all contribute to the same semantic topic. Besides semantic boxes, we also define global boxes, which correspond to the title, header, footer, date, and author of the document. The global boxes are semantically related to all the elements in the page, while the semantic boxes are more local and can be independent of each other. Figure 1 (right) shows four semantic boxes and three global boxes. Introducing global boxes as well as semantic boxes allows us to represent the page’s semantic dependency structure more accurately.
As we are the first to define semantic document layout analysis in this context, we developed our training and evaluation dataset through a rigorous annotation process. This labor-intensive but crucial step was fundamental to achieving a high-quality model. We collaborated with a specialized data annotation company in Tokyo, engaging six expert Japanese annotators, in a reasonable payment and contract. We asked them to locate the semantic boxes and global boxes in the document images and label them with the corresponding types. The box types we defined are: semantic_box, title, header, footer, date, and author.
Source documents for annotation were collected from Common Crawl PDFs Turski et al. (2023). To ensure diversity in our training data, we first performed clustering on the CC-PDF corpus using image embeddings with MiniCPM-Visual-Embedding Rhapsody Group (2024) model. We then selected a balanced number of candidates from each cluster for annotation, creating a representative dataset that encompasses a wide range of document styles, formats, and domain-specific layouts. We finally labeled 24,577 images and used 22,577, 1,000, and 1,000 samples for training, validation, and test sets. Appendix A provides examples of our dataset.
3.1.2 Object Detection Model Finetuning
There are two popular families of object detection architectures: the CNN-based (YOLO series Redmon et al. (2016); Khanam and Hussain (2024)) and the transformer-based (DETR series Carion et al. (2020); Zhao et al. (2024a)). We tried to fine-tune YOLO11-X (56.9M parameters) [^1] https://docs.ultralytics.com/models/yolo11/ with 5 and RT-DETR-X (67.4M parameters) [^2] https://docs.ultralytics.com/models/rtdetr/ models with our semantic layout annotation data supported on the Ultralytics framework (under New AGPL-3.0 License) to develop our SCAN models.
To evaluate the fine-tuned models in terms of bounding box granularity and precision, we developed a matching-based IoU (Intersection over Union) metric. This metric first uses the Hungarian algorithm to perform bipartite matching between predicted and ground-truth bounding boxes. And then, it calculates an average IoU over the matched bounding box pairs and unmatched bounding boxes. For unmatched bounding boxes, we assigned an IoU of 0 to penalize excessive or insufficient predictions. Table 1 shows the two models’ IoU scores on the validation set of our dataset. Because each predicted bounding box has its confidence score, we varied the confidence threshold to select the optimal predicted bounding box set. With the optimal confidence threshold, the performance of the YOLO11-X and RT-DETR-X finetuned models is similar, while the RT-DETR-X-based model is slightly better. We primarily use RT-DETR-X hereafter.
Model Conf. threshold IoU
YOLO11-X 0.2 53.7
0.3 58.0
0.4 55.8
0.5 49.7
RT-DETR-X 0.2 48.9
0.3 54.8
0.4 58.4
0.5 59.6
Table 1: The Intersection over Union (IoU) scores of fine-tuned object detection models. We varied the confidence threshold for predicted bounding boxes from 0.2 to 0.5.
3.2 Predicted Box Selection
Object detection models typically output many bounding boxes along with their corresponding confidence scores. We then select some of the bounding boxes for the final prediction. The simplest selection approach would be to establish a fixed confidence threshold. However, this strategy proves suboptimal for document layout analysis due to document-specific strict constraints:
  • Predicted bounding boxes should cover all textual and visual elements to minimize information loss
  • Predicted bounding boxes should not overlap with each other to minimize information redundancy
This challenge distinguishes document layout analysis from traditional object detection tasks. While conventional object detection aims to identify and localize specific objects with sufficient accuracy, document layout analysis must ensure mutually exclusive and collectively exhaustive predictions.
To address this, we developed an adaptive selection strategy based on two metrics: coverage ratio and non-overlap ratio. The coverage ratio $R_{c}$ measures the proportion of the page area contained within any of the selected boxes:
$R_{c}=\frac{A\left(\bigcup_{i=1}^{n}B_{i}\right)}{A(P)},$ (1)
where $B_{i}$ represents the region of selected box $i$, $A(B)$ represents the area of region $B$, $A(P)$ denotes the whole page area, and $n$ is the number of selected boxes. The non-overlap ratio $R_{o}$ quantifies the degree of redundancy between selected boxes:
$R_{o}=1-\frac{A\left(\bigcup_{1\leq i<j\leq n}(B_{i}\cap B_{j})\right)}{A(P)}.$ (2)
We then define a weighted sum score $S$ to balance these competing objectives:
$S=\alpha\cdot R_{c}+\beta\cdot R_{o},$ (3)
where $\alpha$ and $\beta$ are weighting parameters controlling the trade-off between coverage and overlap.
Our implementation follows a greedy progressive approach: we first establish a minimum confidence threshold to create a candidate pool of boxes. Starting with an empty selection, we iteratively add boxes in descending order of confidence, calculating the score $S$ at each step. A box is added to the final selection if it increases the overall score $S$; otherwise, the box is skipped. This approach ensures we maximize information coverage while minimizing redundancy. We tuned the parameters $\alpha$, $\beta$, and confidence threshold in terms of the IoU score, which resulted in $\alpha=0.9$, $\beta=0.1$, and a confidence threshold of 0.4. We used fine-tuned RT-DETR-X and the validation split of DocSemNet for the parameter tuning.
In addition, to further refine the selection result, we remove boxes completely covered by another selected box, contributing to less redundancy while keeping coverage. After employing these strategies, we achieved an IoU score of 60.2. Appendix C shows the examples before and after the predicted box selection.
3.3 Post-processing for RAG
We first extract the sub-images of these elements by the coordinates of SCAN’s predicted bounding box result. For textual RAG, we use VLMs to convert these sub-images into text, then combine them according to their original reading order relationships to obtain the page-level result for retrieval. We can use general rule-based reading order methods; in our experiments, we simply sorted the reading order based on the coordinates of the upper-left corners of various elements.
For visual RAG, since we aim to use semantic boxes as new retrieval object units to replace original pages with larger information content, we have three options and ablation in Section 4.4 test the performance of them.
1. all-box: Directly using the blocks segmented after semantic layout analysis as new objects for image retrieval. Note that titles, headers, etc., are treated as independent images.
2. semantic_box-only: Using only the content of semantic_boxes as new objects for image retrieval.
3. concatenation: Using a concatenation method to combine global information with semantic box content. Specifically, we append the sub-images of (title, header, footer, date, and author) to each semantic_box sub-image.
3.4 Cost Discussion for VLM text conversion
When applying SCAN for textual RAG, we replace a single high-resolution document page with small sub-images, but the VLM inference cost is almost unchanged. This is because current mainstream VLMs consider number of tokens based on image pixels. For example, the Qwen-2.5-VL model corresponds 28 $\times$ 28 pixels to a single visual token. SCAN divides images into sub-images, but the total number of pixels remains essentially unchanged, or decreased since SCAN skips some white areas. Therefore, the number of input visual tokens remains roughly the same. Although multiple VLM calls will increase the total text tokens due to repeated text conversion prompts, this increase is negligible since visual tokens dominate the computational cost.
During inference, because of the KV cache mechanism, time efficiency generally increases linearly with the length of output tokens. Therefore, for an expect text conversion result, the cost of SCAN’s multiple inferences with sub-images should be comparable to a single inference with the entire page image. Moreover, SCAN’s segmentation embodies a divide-and-conquer approach that can be parallelized, fully utilizing GPU capabilities when resources permit. Our experiments in Section 4.5 empirically validate this analysis.
4 Experiments
4.1 Datasets and Setting
We use three datasets to evaluate the RAG performance. One English dataset OHR-Bench Zhang et al. (2025) and two Japanese datasets, Allganize Allganize.ai (2024) and our in-house BizMMRAG. Every dataset is for both textual and visual RAG. Table 2 gives a detailed summary of these datasets.
Dataset Domain Lang. #Docs #Pages #QA QA Type
OHR-Bench Zhang et al. (2025) Academic, Administration, Finance, En 1,261 8,561 8,498 Text, Table, Formula,
Law, Manual, Newspaper, Textbook Chart, Reading Order
Allganize Allganize.ai (2024) Finance, Information Technology, Manufacturing, Ja 64 2,167 300 Text, Table, Chart
Public Sector, Retail/Distribution
BizMMRAG Finance, Government, Medical, Ja 42 3,924 146 Text, Table, Chart
Consulting sectors, Wikipedia
Table 2: Statistics of our evaluation datasets.
Retrieval Generation Overall
TXT TAB FOR CHA RO ALL TXT TAB FOR CHA RO ALL TXT TAB FOR CHA RO ALL
Ground Truth 81.2 69.6 74.8 70.3 9.8 70.0 49.4 46.0 34.0 47.0 28.2 43.9 45.0 34.6 28.0 32.9 18.7 36.1
Pipeline-based OCR
MinerU 67.7 48.5 51.1 16.5 5.9 50.1 45.9 39.3 28.6 9.7 29.5 36.7 41.4 28.5 23.0 9.3 17.8 30.0
Marker 75.2 57.8 55.4 19.7 5.9 56.6 44.5 37.8 27.8 10.9 26.2 35.9 40.1 28.1 22.3 10.0 16.2 29.5
OCR-specialized small VLM
GOT-580M 62.1 41.0 48.7 17.4 3.7 45.4 37.5 28.5 24.1 8.5 7.1 27.8 35.3 22.9 20.1 8.2 5.3 24.6
Nougat-350M 59.1 32.7 44.2 11.3 4.4 40.9 36.7 22.9 22.9 6.4 6.9 25.5 33.5 18.4 19.4 5.8 3.6 14.5
SCAN-GOT-580M (ours) 60.4 44.1 42.6 28.5 4.5 45.8 ($+$0.4) 45.4 47.4 27.3 20.3 23.2 36.3 ($+$8.5) 38.4 25.2 20.3 15.2 17.4 28.1 ($+$3.5)
SCAN-Nougat-350M (ours) 61.7 27.9 37.1 10.8 4.3 39.5 (--1.4) 41.9 22.7 21.9 6.4 4.1 27.3 ($+$1.8) 37.3 17.9 17.5 6.4 5.4 23.5 ($+$9.0)
General VLM
InternVL2.5-78B 68.6 57.9 55.6 45.1 2.7 56.2 41.7 41.8 29.0 33.6 3.3 35.8 38.2 31.0 23.3 22.9 3.1 29.6
olmOCR-7B 72.5 58.4 55.4 24.8 5.0 56.6 44.8 40.5 30.4 19.0 8.4 36.0 40.6 30.3 23.7 12.8 7.1 29.6
Qwen2.5-VL-72B 75.1 60.0 60.0 38.2 5.3 59.6 44.3 42.1 31.8 27.0 11.6 37.5 40.6 31.1 26.1 19.0 8.8 31.1
DocLayout-YOLO-Qwen2.5-VL-72B 63.5 12.4 36.0 11.7 5.4 36.3 (--23.3) 41.2 10.2 19.0 7.2 18.2 24.4 (--13.1) 38.4 10.0 16.3 7.7 12.5 22.4 (--8.7)
Beehive-Qwen2.5-VL-72B 73.3 16.0 42.3 16.3 6.1 42.6 (--17.0) 46.6 11.3 21.3 10.2 27.2 28.2 (--9.3) 43 10.9 19.5 8.6 16.6 25.3 (--5.8)
SCAN-Qwen2.5-VL-72B (ours) 75.2 56.8 56.6 36.8 6.2 58.3 (--1.3) 47.4 41.7 30.2 23.9 28.3 39.5 ($+$2.0) 43.3 31.3 25.2 17.9 17.5 32.8 ($+$1.7)
Table 3: Textual RAG results on OHR-Bench. Comparison of various OCR methods across different evaluation metrics. TXT, TAB, FOR, CHA, RO, and ALL represent text, table, formula, chart, reading order, and their average, respectively. The RO category includes questions that require identifying the correct reading order to associate information from separate paragraphs.
When evaluating OHR-Bench for textual RAG, we follow the same setting as in the paper Zhang et al. (2025). We use BGE-m3 Chen et al. (2024a) and BM25 as retrieval models, and meta-llama/Llama-3.1-8B-Instruct and Qwen/Qwen2-7B-Instruct as answer models to generate answers according to the retrieved top-2 results. We evaluate three metrics: (a) retrieval, which calculates LCS (Longest Common Subsequence) to measure evidence inclusion in retrieved content; (b) generation, which measures the F1-score of QA when provided with the ground truth page; and (c) overall, which calculates the F1-score of QA for the end-to-end RAG pipeline. The F1-score is calculated using precision and recall of common tokens between the generated result and the ground truth. The final scores will be the average score across the combination of two retrieval models and two answer models. On the other hand, we also apply OHR-Bench to visual RAG, which uses ColQwen2-v1.0 Faysse et al. (2024) as an image retrieval model with top-5 retrieval, and use Qwen/Qwen2.5-VL-7B as an answer model.
For the Japanese datasets BizMMRAG and Allganize, we employ intfloat/multilingual-e5-large as a textual retrieval model, and ColQwen2-v1.0 as an image retrieval model. For answer generation, we employ GPT-4o (gpt-4o-2024-08-06) in textual RAG and Qwen2.5-VL-7B in visual RAG. For each query, the top-5 retrieved results are used. For evaluation, we adopt the LLM-as-a-judge framework Gu et al. (2025), using GPT-4o with a temperature of 0.3 to assign an integer score from 1 to 5 to each generated answer. Answers receiving a score greater than 4 are considered correct (assigned a value of 1), while others are considered incorrect (assigned a value of 0). Final accuracy is computed based on these binary scores.
We test all three post-processing options in visual RAG and report the best result.
BizMMRAG Allganize
TXT TAB CHA ALL TXT TAB CHA ALL
Textual RAG
LlamaParse-Premium 83.3 59.1 69.1 70.5 80.1 59.2 72.0 70.7
Qwen2.5-VL-72B 85.0 52.3 69.5 68.7 84.5 68.4 62.2 71.7
DocLayout-YOLO-Qwen2.5-VL-72B 61.7 25.0 28.6 38.4 (--30.3) 49.3 21.1 23.2 31.2 (--40.5)
Beehive-Qwen2.5-VL-72B 70.0 29.6 21.4 40.3 (--28.4) 61.3 40.8 26.8 43.0 (--28.7)
SCAN-Qwen2.5-VL-72B (ours) 81.0 70.5 73.8 75.3 ($+$6.6) 86.6 79.0 68.3 78.0 ($+$6.3)
Visual RAG
Qwen2.5-VL-7B 71.7 56.8 57.1 58.9 75.9 71.1 62.5 69.9
SCAN-Qwen2.5-VL-7B (ours) 75.0 63.6 57.1 65.3 ($+$6.4) 79.4 64.5 76.3 73.4 ($+$3.5)
Table 4: Textual and visual RAG results on Japanese datasets: BizMMRAG and Allganize.
TXT TAB FOR CHA RO ALL
Qwen2.5-VL-7B 84.0 68.6 71.5 58.7 67.9 70.2
SCAN-Qwen2.5-VL-7B (ours) 86.1 68.6 73.2 61.7 87.8 75.5 ($+$5.3)
Table 5: Visual RAG results on OHR-Bench.
4.2 Textual RAG Evaluation Results
OHR-Bench. Table 3 presents the comprehensive evaluation results of our SCAN method applied to various VLMs for text conversion in textual RAG. Among conventional approaches, general VLMs demonstrate superior performance, followed by pipeline-based OCR methods, and the last are OCR-specialized small VLMs.
Our SCAN model can improve the performance of VLM-based text conversions. Even the strong baseline of Qwen2.5-VL-72B model achieves an impressive overall score of 31.1% (the ground truth is 36.1%), applying SCAN to Qwen2.5-VL-72B further improves the performance to 32.8%. The performance gains are larger when applying SCAN to OCR-specialized small VLMs. With GOT and Nougat, our SCAN improvements are 3.5% and 9.0%, respectively, enabling these smaller models to achieve competitive performance comparable to much larger-sized VLMs. This finding has important implications for deployment scenarios with computational constraints, suggesting that our semantic layout analysis approach can help bridge the efficiency-performance gap. The generation results exhibit similar improvement patterns. The results also indicate that SCAN’s enhancements for structured content elements such as reading order (RO), tables (TAB) and formulas (FOR) become increasingly significant. This suggests that the semantic segmentation approach is particularly valuable for preserving the relationships between elements that have spatial dependencies.
On the other hand, applying SCAN slightly degrades the retrieval performance for some models. This is because retrieval represents a relatively simpler task within the RAG pipeline, primarily requiring the correct identification of keywords rather than comprehensive document understanding. In contrast, the subsequent question-answering stage demands precise and complete conversion of document content into text, where SCAN’s semantic layout analysis proves particularly advantageous.
We can also see that conventional fine-grained document analysis methods substantially degrade overall performance, with DocLayout-YOLO Zhao et al. (2024b) reducing retrieval performance by 23.3% and overall performance by 8.3%. Beehive Auer et al. (2024) has the same trends that reduce 5.8% overall. These degradations are particularly severe for structured content types such as tables and charts. This proves that conventional layout analysis methods typically segment documents into small atomic regions, which also break the structure of documents. In contrast, our semantic box approach preserves the integrity of semantically coherent regions, maintains their holistic structure while still providing the organizational benefits of layout analysis. This preservation of semantic unity enables VLMs to process each region with full contextual awareness, leading to more accurate text conversion and, ultimately, superior RAG performance.
BizMMRAG and Allganize. The upper section of Table 4 presents the textual RAG performance for Japanese document datasets. The results have the same trends as the English OHR-Bench evaluation, demonstrating that our SCAN methodology yields substantial improvements over general VLMs. Specifically, on the BizMMRAG dataset, our SCAN-enhanced approach demonstrates notable 6.6% improvements compared to the baseline Qwen2.5-VL-72B model: text accuracy decreased by 4.0%, while table increased by 18.2% and chart increased by 4.3%. We observe similar trends for the Allganize dataset. It is particularly noteworthy that our SCAN-enhanced method outperforms LlamaParse-Premium LlamaParse.AI (2025), a commercial product offered by LlamaIndex, specifically designed for document conversion. This comparison against a commercial solution validates the practical utility and competitive advantage of our research contribution. In addition, the result shows the same trends: using fine-grained document analysis, DocLayout-YOLO, and Beehive heavily degrades overall performance.
OHR-Bench BizMMRAG Allganize
semantic_box-only 75.1 65.3 72.1
concatenation 73.4 65.3 73.4
all-box 75.5 58.9 71.7
Table 6: Ablation study to compare the different post-processing strategies in visual RAG.
Setting # Input tokens # Output tokens Time (s)
Single-page 11,295 (prompt=255) 1,005 71.3
Multi-chunks 8,079 (prompt=1,071) 1,092 72.7
Table 7: Token numbers, processing time comparison of VLM text conversion.
4.3 Visual RAG Evaluation Results
OHR-Bench. Table 5 presents the result of OHR-Bench in visual RAG. When applying our SCAN approach to divide original pages into semantic chunks and performing visual RAG on these chunks, we observed an overall improvement of 5.3% compared to processing entire pages. While the Table task has the same performance as the baseline, we can see modest improvements in other categories, along with a significant enhancement in the Reading Order task. Recall that Reading Order problems require examining different paragraphs and articles to summarize answers. This demonstrates that dividing a page image into independent semantic chunks can reduce noise in reading order-related problems, enabling the model to provide more accurate responses.
BizMMRAG and Allganize. The lower section of Table 4 presents the visual RAG results for Japanese document datasets. The findings demonstrate that on both the BizMMRAG and Allganize benchmarks, our SCAN methodology exhibits substantial accuracy improvements. Specifically, SCAN improves 6.4% on BizMMRAG and 3.5% on Allganize. This result also shows that our SCAN approach enables the VLM to achieve significantly enhanced performance in Japanese VQA.
4.4 Ablation study of post-processing
Table 6 shows visual RAG results with different post-processing options. all-box performed the on OHR-Bench, while for BizMMRAG and Allganize, concatenation yielded better scores. Upon examining the OHR-Bench data, we found that the questions are generally unrelated to titles, headers, and other global boxes. This could explain why the all-box approach (which indexes different elements separately) and the semantic_box-only approach (which directly filters out title elements) performed relatively well. Conversely, we observed that queries in BizMMRAG and Allganize typically required consideration of both global information and main content, most likely causing the superior performance of the concatenation method for these datasets. For practical applications, we recommend experimenting with different approaches to select the best post-processing option.
4.5 Cost Comparison of VLM Text Conversion
Table 7 compares the processing time of VLM-based text conversion between single-page processing and multiple semantic chunks processing. [^3] We used Qwen-2.5-VL-72B locally for this experiment. We can see that the average processing time for semantic chunks (our SCAN approach) is similar to the single-page processing time. The majority of input tokens are vision tokens, compared to the text tokens from conversion prompts. While the total input tokens for multiple chunks are fewer, the output tokens are more. This shows that our SCAN approach can reduce token-based costs (especially favorable for pay-as-token-usage API models) while extracting more textual information, all within approximately the same processing time.
5 Conclusion
We presented SCAN, a semantic document layout analysis approach for modern textual and visual RAG systems. By introducing coarse-grained semantic segmentation that preserves topical coherence, SCAN effectively reduces the information processing burden on VLMs while maintaining semantic integrity across document components. To develop SCAN, we labeled more than 24k document images with semantic layouts and trained a robust semantic layout analysis model. Our comprehensive evaluation across multiple datasets, languages, and document types demonstrates SCAN’s ability to enhance 1.7%–9.0% for textual RAG and 3.5%–6.4% for visual RAG. In addition, SCAN achieves these improvements while reducing token usage costs and maintaining the same processing time.
Ethical Statement
In this work, we study semantic document layout analysis for RAG. To the best of our knowledge, there is no negative societal impact in this research. All our training data consists of publicly available PDFs from the internet, which similarly presents no ethical concerns. Our SCAN model aims to improve information extraction without introducing biases in the underlying content. We believe that improved document analysis can enhance accessibility of information for users across different languages and document formats. While our system improves RAG capabilities, users should still be mindful of the general limitations of AI systems when relying on generated answers.
We used Claude-3.7-Sonnet to polish the writing of the paper. We are responsible for all the materials presented in this work.
Limitations
While our SCAN approach offers significant advantages, we acknowledge several limitations that present opportunities for future research:
1. Our SCAN model operates based on spatial image layout. In certain documents where contents that should logically form a single semantic chunk is physically separated in space and do not fit in a rectangular, our current model cannot yet establish these connections. This limitation could potentially be addressed through an additional trainable reading order model coupled with a semantic_box merging mechanism.
2. Our current model was trained primarily on Japanese data. While experiments demonstrate improvements on English benchmarks as well, this may not represent the optimal model for all languages. Japanese documents have unique layout characteristics, such as vertical writing and right-to-left orientation, which differ from English conventions. Further analysis and exploration are needed, and future work could involve annotating purely English data to investigate whether higher performance could be achieved for English RAG applications.
3. SCAN’s semantic layout was designed for dense, content-rich document RAG. For simpler pages, designing an adaptive approach that intelligently decides whether to apply semantic layout analysis or process the page as a single unit could provide better generalizability in future iterations.
Appendix A Examples from Our Semantic Layout Dataset
Figure 3 shows some examples of our semantic layout dataset. It contains diverse document pages, including research papers, administrative reports, user manuals, slides, flyers, and more. The dataset is annotated with the semantic_box class and the five global box classes: title, header, footer, date, and author.
Appendix B Training Details of Our Object Detection Models
We fine-tuned two off-the-shelf object detection models: YOLO11-X and RT-DETR-X. We followed the default settings provided by the Ultralytics framework (version 8.3.28) [^4] https://github.com/ultralytics/ultralytics/releases/tag/v8.3.28 mostly, but we explicitly set or tuned some important hyper-parameters as shown in Table 8. For the YOLO11-X fine-tuning, we used 4 NVIDIA L40 GPUs (48GB), which took about 4 hours to finish 30 epochs. For the RT-DETR-X fine-tuning, we used 8 NVIDIA A100 GPUs (80GB), which took about 16 hours to finish 120 epochs.
Appendix C Examples of Before and After the Predicted Box Selection
Figure 46 exemplify our predicted box selection strategy for the semantic document layout analysis.
Appendix D Experimental Details
D.1 Environment
For the text conversion with general VLMs, we used 8 NVIDIA L40S (48GB) GPUs with INTEL(R) XEON(R) GOLD 6548N CPU, and the details are as follows.
python 3.12
vllm==0.7.3 (V0 version)
torch==2.5.1
torchaudio==2.5.1
torchvision==0.20.1
transformers==4.49.0
ultralytics==8.3.28
vLLM settings
- temperature: 0.3
- top_p: 0.95
- max_tokens: 8192
- repetition_penalty: 1.1
- tensor_parallel==4
D.2 Prompts for the VLM Text Conversion
You are a powerful OCR assistant tasked with converting PDF images to the Markdown format. You MUST obey the following criteria:
1. Plain text processing:
- Accurately recognize all text content in the PDF image without guessing or inferring.
- Precisely recognize all text in the PDF image without making assumptions in the Markdown format.
- Maintain the original document structure, including headings, paragraphs, lists, etc.
2. Formula Processing:
- Convert all formulas to LaTeX.
- Enclose inline formulas with \$ \$. For example: This is an inline formula \$ E = mc^2 \$.
[…]
3. Table Processing:
- Convert all tables to LaTeX format.
[…]
4. Chart Processing:
- Convert all Charts to LaTeX format.
[…]
5. Figure Handling:
- Ignore figures from the PDF image; do not describe or convert images.
6. Output Format:
- Ensure the Markdown output has a clear structure with appropriate line breaks.
- Maintain the original layout and format as closely as possible.
Please strictly follow these guidelines to ensure accuracy and consistency in the conversion. Your task is to accurately convert the content of the PDF image using these format requirements without adding any extra explanations or comments.
D.3 Prompts for LLM-as-a-judge
System:
You are an expert evaluation system for a question answering chatbot.
You are given the following information:
- a user query and reference answer
- a generated answer
You may also be given a reference answer to use for reference in your evaluation.
Your job is to judge the relevance and correctness of the generated answer.
Output a single score that represents a holistic evaluation.
You must return your response in a line with only the score.
Do not return answers in any other format.
On a separate line provide your reasoning for the score as well.
Follow these guidelines for scoring:
- Your score has to be between 1 and 5, where 1 is the worst and 5 is the best.
- Your output format should be in JSON with fields "reason" and "score" shown below.
- If the generated answer is not relevant to the user query, you should give a score of 1.
- If the generated answer is relevant but contains mistakes, you should give a score between 2 and 3.
- If the generated answer is relevant and fully correct, you should give a score between 4 and 5.
Example Response in JSON format:
"reason": "The generated answer has the exact same metrics as the reference answer, but it is not as concise.",
"score": "4.0"
User:
## User Query
{query}
## Reference Answer
{reference_answer}
## Generated Answer
{generated_answer}
Figure 3: Examples of our semantic layout dataset.
Hyper-parameter YOLO11-X RT-DETR-X
Batch size {8, 16, 32, 64} {8, 16, 32, 64}
Learning rate {1e-4, 5e-4, 1e-3, 5e-3} {5e-5, 1e-4, 5e-4, 1e-3, 5e-3}
Max training epochs {30, 40, 80, 120} {80, 120, 160}
Weight decay {5e-5, 1e-4, 5e-4} {1e-5, 1e-4, 1e-3}
Warmup epochs {5, 10} {5, 10}
Image size 1024
Dropout 0.0
Optimizer AdamW
Learning rate scheduler cos_lr
Table 8: Hyper-parameters used for fine-tuning object detection models. We tuned the hyper-parameters in the brackets in terms of the mean average precision (mAP) on the validation set. The bold values are the best hyper-parameters for each model.
Figure 4: Semantic document layout analysis result before (left) and after (right) our predicted box selection strategy (1/3).
Figure 5: Semantic document layout analysis result before (left) and after (right) our predicted box selection strategy (2/3).
Figure 6: Semantic document layout analysis result before (left) and after (right) our predicted box selection strategy (3/3).